OpenStudy (anonymous):

iamabear In the TV show "Breaking Bad" the characters attempt to use HF acid to dissolve guns (among other things). Here we consider instead dissolving guns (which we will assume are pure iron) with sulfuric acid. Complete the balanced reaction for reacting iron in dilute sulfuric acid to form aqueous FeSO4. Do not worry about formatting subscripts (i.e. O2 to represent diatomic oxygen gas is fine). Fe + H2SO4 → FeSO4 + __H2(g)___

OpenStudy (anonymous):

It is apparent that they must use an outside source of electrical power to drive the dissolution of the iron. They use a small current so that the dissolution proceeds with the minimum voltage required. Assume standard values for the electrochemical potentials. How much electrical energy supplied this way is thus required to dissolve an additional 1 kg of iron? Give your answer in kJ.

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

okay, so let's think about it. The reaction is \(Fe \rightarrow 2 e + Fe^{2+}\) and \(2H^+ + 2e \rightarrow H_2\) so if we use \(E_{cell}=E^{red}_{red}-E^{oxd}_{red}\), \(E_{Cell}=0-(-0.44)\) i think we need to use: \(\Delta G=nFE_{cell} \) \(\Delta G=(2)*(9648.5 C/mol)*(0.44~J/C)=8490.68~J/mol\) 1000g/55.8=17.9 moles of iron 8490.68 J/mol*17.9 moles of iron= 151983.172 J = 152 kJ it seems like a lot, though

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

152 = wrong answer

OpenStudy (aaronq):

so where do you think i went wrong?

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

where is the 2kg?

OpenStudy (anonymous):

@aaronq is that 2kg?

OpenStudy (anonymous):

maybe the question is not a good question.

OpenStudy (aaronq):

i'm not sure if it's supposed to be 1 kg or 2 kg. It says "to dissolve an additional kg".

OpenStudy (aaronq):

you could try multiplying that number by 2

OpenStudy (anonymous):

but 152 is wrong....lets try 2x

OpenStudy (anonymous):

its 304 and wrong still

OpenStudy (anonymous):

your raw answer: 151.983172 is also wrong...lol

OpenStudy (anonymous):

151.983172 and 152 and 304 are all wrong answers....i duno man....so hard

OpenStudy (aaronq):

do you have any of these types of problems in your notes?

OpenStudy (anonymous):

no lol that why im stuck

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

is electrical energy this? http://hyperphysics.phy-astr.gsu.edu/hbase/electric/elepe.html#c1

OpenStudy (anonymous):

the question is under engineering and solar power mechanisms

OpenStudy (aaronq):

oh wait, i wrote faradays constant wrong, it's 96,485 C/mol so: 0.44*2*96485=8490.68 J/mol 8490.68 J/mol*17.9 mol=1519831.72 J = 1520 kJ

OpenStudy (anonymous):

doesn't work still

OpenStudy (anonymous):

wait let me try raw

OpenStudy (anonymous):

let me try without rounded

OpenStudy (aaronq):

this is electrochemistry, related, but not the same as the link you posted.

OpenStudy (anonymous):

one moment

OpenStudy (anonymous):

hm none of teh answer work

OpenStudy (anonymous):

@aaronq none of the answer works both rounded and non-rounded, but it only accepts raw numbers un-rounded

OpenStudy (anonymous):

@aaronq lets just give up ok?

OpenStudy (aaronq):

hmm sure. I don't think i'm making a mistake, but leave it open. Maybe someone will come and take a look and spot where we went wrong.

OpenStudy (anonymous):

you are right but i dont think it is in the right format or u might have missed something the question wants you to do

OpenStudy (anonymous):

i do not doubt that your answer is wrong but the system didnt accept it.

OpenStudy (aaronq):

do you think it's asking for the energy for 2 kg?

OpenStudy (anonymous):

we ar just missing something...could be another number factor or somtin i do not know

OpenStudy (anonymous):

multiply it by 2?

OpenStudy (anonymous):

2279.74758 doesn't work

OpenStudy (anonymous):

gives a X wrong

OpenStudy (aaronq):

Its weird that they said "an additional 1 kg". "Additional" implies that there is more than just the 1 kg. Let's leave it. Hopefully someone spots the error.

OpenStudy (anonymous):

maybe we take a break im tired lol

OpenStudy (aaronq):

haha sounds good.